OpenStudy (anonymous):

Easy Question!! The side of a square is measured to be 16 ft with a possible error of ±0.1 ft. Use differentials to estimate the error in the calculated area. Include units in your answer

OpenStudy (anonymous):

x=16 \[dx=\pm0.1\] \[Area ~A=x^2\] \[\frac{ dA }{ dx }=2x\] dA=2x~dx \[dA=2*16*\left( \pm0.1 \right)=?\]

OpenStudy (anonymous):

well 32 plus or minus 0.1 right?

OpenStudy (anonymous):

\[\pm~3.2\]

OpenStudy (anonymous):

oh okay so you multiply the .1 by the 32 , i see

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Thank you very much !! :D

OpenStudy (anonymous):

yw